5. Solve the following system by addition. If a unique solution does not exist, state
whether the system is inconsistent or dependent.
2x + 3y = 1
5x + 3y = 16


Answer by checkley71(8403)

2X+3Y=1
5X+3Y=16 SUBTRACTING EQ#2 FROM EQ#1 WE GET
----------
-3X=-15 OPR 3X=15 OR X=15/3 OR X=5 THEN 2*5+3Y=1 OR 10+3Y=1 OR 3Y=-9 OR Y=-3